Question 2 options:

Given the value for the universal gravitational constant, the average value for the acceleration due to gravity at the surface of Earth, and the radius of Earth, Earth’s mass is ____x1024kg


1
Expert's answer
2021-07-28T03:56:02-0400

We have, acceleration due to gravity,


"g = \frac {GM} {R2}"


Where, G = Gravitational constant


= "6.67\u00d710^{-11} N.m^2\/kg^2,"


R = radius of the Earth = "6.4\u00d710^6 m" ,


g = acceleration due to gravity on Earth surface


= 9.8 m/s2 and M is the mass of Earth, then, we can write,


"9.8 = 6.67×10^{-11} ×\frac {M} {(6.4×10^6)^2}"


=> "9.8 = M\u00d70.16284\u00d710^{-23}"


So, mass of Earth = "\frac {9.8} {(0.16284×10^{-23} ) kg}"


= "5.97\u00d710^{24} kg"